OpenStudy (anonymous):

determine the Fourier series representation for: f(x)=x;0≤x≤π 1; π

OpenStudy (anonymous):

\[a _{n}=(1/2)\int\limits_{0}^{\pi}xcos (n \pi x / 2) dx +(1/2)\int\limits_{\pi}^{2\pi} \cos (n \pi x/2)dx\] try to integrate that for n not equal to zero;

OpenStudy (anonymous):

and for n=0 use \[a _{0=}(1/2)\int\limits_{0}^{\pi}xdx +(1/2)\int\limits_{\pi}^{2\pi}dx\]

OpenStudy (anonymous):

\[and for] n \neq0 ] use\] \[b _{n}=(1/2)\int\limits_{0}^{\pi} xsin(n \pi x/2)+(1/2)\int\limits_{\pi}^{2\pi}\sin(n \pi x/2)\]

OpenStudy (anonymous):

and substitute them in the fourier series form \[f(x)=(1/2) a _{0} + \sum_{n=1}^{\infty}( a _{n} \cos(n \pi x/2) +b _{n}\sin (n \pi x/2))\]

OpenStudy (anonymous):

have fun and good luck fazreenlyana

OpenStudy (anonymous):

i still confused when to used the even and odd function.. n if the question does not mention about the number of N, what we should consider?? can you help me to make me clear about this situation..